# Replay-Free Continual Low-Rank Adaptation with Dynamic Memory

### v3

@@ 4.2 Residual Adapter
As the feature subspaces represented by $\boldsymbol{\Phi}_{t}^{k}$ and $\boldsymbol{\Phi}_{t+1}^{v}$ expand with the accumulation of learned tasks, the majority of the components in the updates, $\Delta\mathbf{O}_{t+1}^{k}$ and $\Delta\mathbf{O}_{t+1}^{v}$, are progressively subtracted, as detailed in (7). Consequently, the update magnitudes approach zero, resulting in diminished performance during fine-tuning on new tasks. To address this issue, we introduce a residual adapter $\mathbf{R}_{t+1}$ (initialized as $\mathbf{0}$) in parallel with $\mathbf{O}_{t+1}^{v}$, providing extra capacity for new tasks and maintaining a balance between stability and plasticity.
@@
$\boldsymbol{\Psi}_{t}:=\boldsymbol{\Phi}_{t}^{v}-\boldsymbol{\Phi}_{t-1}^{v}\subseteq\mathbb{R}^{d},$ (8)
where $\boldsymbol{\Phi}_{t}^{v}$ and $\boldsymbol{\Phi}_{t-1}^{v}$ are obtained after fine-tuning on the $t$ and t1t-1 tasks. It is worth noting that subspace $\mathcal{R}_{t}$ and $\mathcal{R}_{t+1}$ are specific to their corresponding tasks $t$ and $t+1$, respectively, as $\boldsymbol{\Psi}_{t+1}\cap\boldsymbol{\Psi}_{t}=\emptyset$. Specifically, the subspace $\mathcal{R}_{t}$ indicates the residual knowledge extracted from the most recent task, providing supplementary bases to enlarge the optimization subspace. With these extracted bases, we are able to project the updates $\Delta\mathbf{R}_{t+1}$ into the subspace $\mathcal{R}_{t}$ by
@@
When we conduct fine-tuning on the $(t+1)$-task, the value matrix $\mathbf{V}^{(l)}$ in the $l$-th layer given activations $\mathbf{h}^{(l)}$ can be found as
@@
where $\mathbf{W}_{0}^{v}$ is the pre-trained weights, and $\mathbf{V}_{o}^{(l)}:=\mathbf{a}^{(l)}\left(\mathbf{W}_{0}^{v}+\mathbf{O}_{t+1}^{v}\right)$, $\mathbf{V}_{r}^{(l)}:=\mathbf{a}^{(l)}\mathbf{R}_{t+1}$.
@@ 4.3 Dynamic Memory
As previously mentioned, the residual adapter $\mathbf{R}_{t+1}$ is updated within the subspace $\mathcal{R}_{t}$, a subset of the feature subspace spanned by $\boldsymbol{\Phi}_{t}$ extracted from the $t$-th task. Consequently, the fine-tuning process may deteriorate the performance on prior tasks. To mitigate this issue, we introduce a dynamic memory mechanism that adjusts value of $\mathbf{V}_{r}^{(l)}$, which is the output of $\mathbf{R}_{t+1}$ during inference on test data according to
$\hat{\mathbf{V}}^{(l)}=\mathbf{V}_{o}^{(l)}+\mathbf{a}^{(l)}\boldsymbol{\Omega}_{t+1}^{T}\boldsymbol{\Omega}_{t+1}\mathbf{R}_{t+1}=\mathbf{V}_{o}^{(l)}+\hat{\mathbf{V}}_{r}^{(l)},$ (11)
where $\boldsymbol{\Omega}_{t+1}^{\top}\boldsymbol{\Omega}_{t+1}$ is computed according to the input activation signal $\mathbf{a}^{(l)}$. Specifically, the attention score $\mathbf{S}^{(l)}\propto\text{softmax}(\frac{\mathbf{Q}^{(l)}(\mathbf{K}^{(l)})^{\top}}{\sqrt{d}})$, computed by applying $\mathbf{a}^{(l)}$ to the query and key weight matrices, reflects the relevance between the input test sample and the task associated with the extracted bases used to update the residual adapter $\mathbf{R}_{t+1}$. We utilize the matrices $\boldsymbol{\Psi}_{\tau}\in\mathbb{R}^{r_{\tau}\times d}$ ($\tau\leq t+1$), stored in memory, to multiply the first row of the attention score, $\mathbf{v}^{(l)}:=\mathbf{S}_{1}^{(l)}$, as follows:
 $\text{ if }r_{\tau}\not=0,\omega_{\tau}=\frac{\left\|\boldsymbol{\Psi}_{\tau}\cdot\mathbf{v}^{(l)}\right\|}{r_{\tau}\left\|\mathbf{v}^{(l)}\right\|},\text{ otherwise, }\omega_{\tau}=0,$ (12)
where $r_{\tau}$ indicates the rank of $\boldsymbol{\Psi}_{\tau}$. Since the task-specific residual bases remain independent across different tasks, the cosine similarity between feature vectors $\mathbf{v}^{(l)}$ extracted from input test samples and the stored residual bases $\boldsymbol{\Psi}_{\tau}$ for each task can be used as a scaling factor for the corresponding components in the outputs of the residual adapter. This method assigns lower weights to components irrelevant to the current test samples, while components with high relevance to the samples are given proportionally higher weights. The resulting matrix, $\boldsymbol{\Omega}_{t+1}$, is obtained by:
$\boldsymbol{\Omega}_{t+1}=\left[\begin{array}[]{cccc}\boldsymbol{\Sigma}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{2}\\ \vdots&&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\boldsymbol{\Sigma}_{t+1}\\ \end{array}\right]\left[\begin{array}[]{c}\boldsymbol{\Psi}_{1}\\ \boldsymbol{\Psi}_{2}\\ \vdots\\ \boldsymbol{\Psi}_{t+1}\\ \end{array}\right]\in\mathbb{R}^{r^{\prime}\times d},$ (13)
where $r^{\prime}=\sum_{\tau=1}^{t+1}r_{\tau}$, $\boldsymbol{\Sigma}_{\tau}\in\mathbb{R}^{r_{\tau}\times r_{\tau}}$ is a diagonal matrix with the identical value $\omega_{\tau}^{1/2}$.
@@ 4.4 Task Identification with Confidence
As mentioned earlier, we sample $m$ training data points to extract orthogonal bases after completing the $t$-th task. During this process, we obtain the average feature vector $\bar{\mathbf{v}}^{(L)}$ forwarded to the final attention block from each task and compute the similarity vector $\boldsymbol{\pi}_{t}=\{\omega_{\tau}\}_{\tau=1}^{t}$ based on (12). Therefore, we obtain a set $\boldsymbol{\Pi}_{t}=\{\boldsymbol{\pi}_{1},\dots,\boldsymbol{\pi}_{t}\}$ that can be used to distinguish the task identity during inference. Specifically, let $\boldsymbol{\pi}^{\star}$ denote the similarity vector computed from the input test sample, then we can predict the task identity by
$\hat{k}=\text{arg}\max_{\tau}g(\boldsymbol{\pi}_{\tau},\boldsymbol{\pi}^{\star}):=\frac{\left\|\boldsymbol{\pi}_{\tau}\cdot\boldsymbol{\pi}^{\star}\right\|}{\left\|\boldsymbol{\pi}_{\tau}\right\|\cdot\left\|\boldsymbol{\pi}^{\star}\right\|},$ (14)
$\hat{\delta}=\lambda\left(g(\boldsymbol{\pi}_{\hat{k}},\boldsymbol{\pi}^{\star})-\max_{\tau\not=\hat{k}}g(\boldsymbol{\pi}_{\tau},\boldsymbol{\pi}^{\star})\right),$ (15)
where $\lambda$ is a scaling factor, $\hat{k}$ denotes the predicted task identity and $\hat{\delta}$ indicates the confidence of this prediction. Moreover, we scale the output logits as

### v4

@@ 4.2 Residual Adapter
As the feature subspaces represented by $\boldsymbol{\Phi}_{t}^{k}$ and $\boldsymbol{\Phi}_{t}^{v}$ expand with the accumulation of learned tasks, the majority of the components in the updates, $\Delta\mathbf{O}_{t+1}^{k}$ and $\Delta\mathbf{O}_{t+1}^{v}$, are progressively subtracted, as detailed in (7). Consequently, the update magnitudes approach zero, resulting in diminished performance during fine-tuning on new tasks. To address this issue, we introduce a residual adapter $\mathbf{R}_{t+1}$ (initialized as $\mathbf{0}$) in parallel with $\mathbf{O}_{t+1}^{v}$, providing extra capacity for new tasks and maintaining a balance between stability and plasticity.
@@
$\boldsymbol{\Psi}_{t}:=\boldsymbol{\Phi}_{t}^{v}\setminus\boldsymbol{\Phi}_{t-1}^{v}\subseteq\mathbb{R}^{d},$ (8)
where $\boldsymbol{\Phi}_{t}^{v}$ and $\boldsymbol{\Phi}_{t-1}^{v}$ are obtained after fine-tuning on the $t$ and t1t-1 tasks. It is worth noting that the subspaces $\mathcal{R}_{t}$ and $\mathcal{R}_{t+1}$ are specific to their corresponding tasks $t$ and $t+1$, respectively, as $\boldsymbol{\Psi}_{t+1}\cap\boldsymbol{\Psi}_{t}=\emptyset$. Specifically, the subspace $\mathcal{R}_{t}$ indicates the residual knowledge extracted from the most recent task, providing supplementary bases to enlarge the optimization subspace. With these extracted bases, we are able to project the updates $\Delta\mathbf{R}_{t+1}$ into the subspace $\mathcal{R}_{t}$ by
@@
When we conduct fine-tuning on the $(t+1)$-th task, the value matrix $\mathbf{V}^{(l)}$ in the $l$-th layer given the input activations $\mathbf{a}^{(l)}$ can be found as
@@
where $\mathbf{W}_{0}^{v}$ denotes the pre-trained weights, and $\mathbf{V}_{o}^{(l)}:=\mathbf{a}^{(l)}\left(\mathbf{W}_{0}^{v}+\mathbf{O}_{t+1}^{v}\right)$, $\mathbf{V}_{r}^{(l)}:=\mathbf{a}^{(l)}\mathbf{R}_{t+1}$.
@@ 4.3 Dynamic Memory
As previously mentioned, the residual adapter $\mathbf{R}_{t+1}$ is updated within the subspace $\mathcal{R}_{t}$, a subset of the feature subspace spanned by $\boldsymbol{\Phi}_{t}$ extracted from the $t$-th task. Consequently, the fine-tuning process may deteriorate the performance on prior tasks. To mitigate this issue, we introduce a dynamic memory mechanism that adjusts the value of $\mathbf{V}_{r}^{(l)}$, which is the output of $\mathbf{R}_{t+1}$ during inference on test data according to
$\hat{\mathbf{V}}^{(l)}=\mathbf{V}_{o}^{(l)}+\mathbf{a}^{(l)}\boldsymbol{\Omega}_{t+1}^{\top}\boldsymbol{\Omega}_{t+1}\mathbf{R}_{t+1}=\mathbf{V}_{o}^{(l)}+\hat{\mathbf{V}}_{r}^{(l)},$ (11)
where $\boldsymbol{\Omega}_{t+1}^{\top}\boldsymbol{\Omega}_{t+1}$ is computed according to the input activation signal $\mathbf{a}^{(l)}$. Specifically, the attention score $\mathbf{S}^{(l)}\propto\text{softmax}\!\left(\frac{\mathbf{Q}^{(l)}(\mathbf{K}^{(l)})^{\top}}{\sqrt{d}}\right)$, computed by applying $\mathbf{a}^{(l)}$ to the query and key weight matrices, reflects the relevance between the input test sample and the task associated with the extracted bases used to update the residual adapter $\mathbf{R}_{t+1}$. We utilize the matrices $\boldsymbol{\Psi}_{\tau}\in\mathbb{R}^{r_{\tau}\times d}$ ($\tau\leq t+1$), stored in memory, to multiply the first row of the attention score, $\mathbf{v}^{(l)}:=\mathbf{S}_{1}^{(l)}$, as follows:
$\text{if }r_{\tau}\neq 0,\;\omega_{\tau}=\frac{\left\|\boldsymbol{\Psi}_{\tau}\mathbf{v}^{(l)}\right\|}{r_{\tau}\left\|\mathbf{v}^{(l)}\right\|},\quad\text{otherwise, }\omega_{\tau}=0,$ (12)
where $r_{\tau}$ indicates the rank of $\boldsymbol{\Psi}_{\tau}$. Since the task-specific residual bases remain independent across different tasks, the cosine similarity between feature vectors $\mathbf{v}^{(l)}$ extracted from input test samples and the stored residual bases $\boldsymbol{\Psi}_{\tau}$ for each task can be used as a scaling factor for the corresponding components in the outputs of the residual adapter. This method assigns lower weights to components irrelevant to the current test samples, while components with high relevance to the samples are given proportionally higher weights. The resulting matrix $\boldsymbol{\Omega}_{t+1}$ is obtained by
$\boldsymbol{\Omega}_{t+1}=\left[\begin{array}[]{cccc}\boldsymbol{\Sigma}_{1}&\mathbf{0}&\cdots&\mathbf{0}\\ \mathbf{0}&\boldsymbol{\Sigma}_{2}&\cdots&\mathbf{0}\\ \vdots&&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\cdots&\boldsymbol{\Sigma}_{t+1}\end{array}\right]\left[\begin{array}[]{c}\boldsymbol{\Psi}_{1}\\ \boldsymbol{\Psi}_{2}\\ \vdots\\ \boldsymbol{\Psi}_{t+1}\end{array}\right]\in\mathbb{R}^{r^{\prime}\times d},$ (13)
where $r^{\prime}=\sum_{\tau=1}^{t+1}r_{\tau}$ and $\boldsymbol{\Sigma}_{\tau}\in\mathbb{R}^{r_{\tau}\times r_{\tau}}$ is a diagonal matrix with identical entries $\omega_{\tau}^{1/2}$.
@@ 4.4 Task Identification with Confidence
As mentioned earlier, we sample $m$ training data points to extract orthogonal bases after completing the $t$-th task. During this process, we obtain the average feature vector $\bar{\mathbf{v}}^{(L)}$ forwarded to the final attention block from each task and compute the similarity vector $\boldsymbol{\pi}_{t}=\{\omega_{\tau}\}_{\tau=1}^{t}$ based on (12). Therefore, we obtain a set $\boldsymbol{\Pi}_{t}=\{\boldsymbol{\pi}_{1},\dots,\boldsymbol{\pi}_{t}\}$ that can be used to distinguish the task identity during inference. Specifically, let $\boldsymbol{\pi}^{\star}$ denote the similarity vector computed from the input test sample; then we can predict the task identity by
$\hat{k}=\arg\max_{\tau}g(\boldsymbol{\pi}_{\tau},\boldsymbol{\pi}^{\star}),$ (14)
$g(\boldsymbol{\pi}_{\tau},\boldsymbol{\pi}^{\star}):=\frac{\left\|\boldsymbol{\pi}_{\tau}\cdot\boldsymbol{\pi}^{\star}\right\|}{\left\|\boldsymbol{\pi}_{\tau}\right\|\cdot\left\|\boldsymbol{\pi}^{\star}\right\|},$ (15)
$\hat{\delta}=\lambda\left(g(\boldsymbol{\pi}_{\hat{k}},\boldsymbol{\pi}^{\star})-\max_{\tau\neq\hat{k}}g(\boldsymbol{\pi}_{\tau},\boldsymbol{\pi}^{\star})\right),$ (16)
where $\lambda$ is a scaling factor, $\hat{k}$ denotes the predicted task identity, and $\hat{\delta}$ indicates the confidence of this prediction. Moreover, we scale the output logits as